How many grams of glucose do you need to dissolve in 500mL of water to make a 2 molar solution


1
Expert's answer
2021-03-18T02:59:53-0400

Glucose has molar Mass of 180.16 g/M.

2M/L solution will require 2 ×180.6 = 360.32 g of glucose in 1L water

500 ml into L = 0.5L

= 0.5 ×361.2 = 180.16 grams of glucose to dissolve 500 ml of water